OpenStudy (anonymous):

Will give medal!! please help :) Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=(x-8)/(x+7) g(x)= (-7x-8)/(x-1)

OpenStudy (anonymous):

So f(g(x)) is basically placing the function g(x) into any variable x in f(x) \[f(g(x))=\frac{(\frac{-7x-8}{x-1})-8}{(\frac{-7x-8}{x-1})+7}\] Now solve and simplify and show that f(g(x)=x Now we do the same thing for g(f(x))

OpenStudy (anonymous):

\[g(f(x))=\frac{-7(\frac{x-8}{x+7})-8}{(\frac{x-8}{x+7})-1}\] And now we simplify and show that g(f(x))=x

OpenStudy (anonymous):

not seeing how that equals x

OpenStudy (anonymous):

Well you need to simplify What have you tried so far?

OpenStudy (anonymous):

well the (x-8)/(x+7) is in the numerator and denominator so I cancelled those out

OpenStudy (anonymous):

which legt -7(-8)/ -1

OpenStudy (anonymous):

left*

OpenStudy (anonymous):

\[(g(x))=\frac{(\frac{-7x-8}{x-1})-8}{(\frac{-7x-8}{x-1})+7}=\frac{(\frac{-7x-8}{x-1})-\frac{8(x-1)}{x-1}}{(\frac{-7x-8}{x-1})+\frac{7(x-1)}{x-1}}=\frac{(\frac{-7x-8-8x+8}{x-1})}{(\frac{-7x-8+7x-7}{x-1})}=\frac{\frac{-15x}{x-1}}{\frac{-15}{x-1}}=\frac{-15x}{-15}=x\]

OpenStudy (anonymous):

Okk the last part unfortunately got cut off but here it is \[=\frac{-15x}{-15}=x\]